) Demonstrate by determining the intercepts of "14x + 7y = 21". Non graphing
Answer by checkley77(12844)

14x + 7y = 21
7y=-14x+21 the y intercept=(0,21)
Setting y=0 & solving for x we get:
7*0=-14x+21
-14x=21
x=21/-14
x=-3/2=-1.5 Thus the x intecept=(-1.5,0)